OpenStudy (anonymous):

I will give anyone that helps me a medal A system of equations is shown below. -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} -3x + 7y = -16&{\color{red}{ \times -3}}\implies &\quad \cancel{9x}-21y=48\\ -9x + 5y = 16&&\cancel{-9x} + 5y = 16\\ \hline\\ &&\qquad\quad -16y=64 \end{array}\) so, what do you think would "y" be if you were to solve for "y" there ?

OpenStudy (anonymous):

-4?

OpenStudy (jdoe0001):

yeap now that you know what "y" is, you can just plug it on either equation, and solve for "x" that's the so-called solving by ELIMINATION, we ELIMINATED the "x" by using a multiple on one and then summing them up

OpenStudy (anonymous):

Oh, thank you c: Was that Part A or B?

OpenStudy (jdoe0001):

well, part A asks to do as shown above part B only says that you have to say, grab the RESULTANT equation from the sum and check if the values are the same as the original equations

OpenStudy (anonymous):

How do you do that? :o

OpenStudy (jdoe0001):

hmmm lemme see if I get .. A.... .seems to ask for a yet another set

OpenStudy (anonymous):

kk :3

OpenStudy (jdoe0001):

.... the wording is a bit jumbled .... but basically is what was done since you ended up with a new equation from the vertical SUM and a multiple of one of the equations

OpenStudy (anonymous):

Ok, but how do I write down the steps? The end part got confusing :P

OpenStudy (anonymous):

Oh, and how do I do Part B?

OpenStudy (jdoe0001):

one could say that your original equations were -3x + 7y = -16 -9x + 5y = 16 and after doing the ELIMINATION we had another one, a new one you can plug into the originals so one could say that a new system of equations could be rewritten like -3x + 7y = -16 or 0x -16y = 64 0x -16y = 64 -9x + 5y = 16

OpenStudy (jdoe0001):

part B means, solve both systems and see that the "x" and "y" on both are the same so the new system per se, has the same solutions as the original system

OpenStudy (anonymous):

Ooooooooooh, I get it :P, Thank you :3

OpenStudy (jdoe0001):

yw